Solve the following differential equation subject to the given initial conditions. 1. dy/dtheta=ysin(theta); y(pi)=3. 2. x^2 dy/dx=y-xy; y(1)=1


1
Expert's answer
2021-07-19T07:43:06-0400

Solution.

1.


"\frac{dy}{d\theta}=y\sin{\theta}, y(π)=3""\frac{dy}{y}=\sin{\theta}d\theta,""\int\frac{dy}{y}=\int\sin{\theta}d\theta,""\ln |y|=-\cos{\theta}+C,"

where C is some constant.

If "y(\u03c0)=3," then "\ln 3=-\cos{π}+C."

From here

"C=\ln 3-1."

We will have

"\ln y=-\cos{\theta}+\ln 3-1," or"\ln y=\ln e^{-\cos{\theta}}+\ln 3-\ln e."

Answer.


"y=\frac{3e^{-\cos{\theta}}}{e}"

2.


"x^2\frac{dy}{dx}=y-xy, y(1)=1""\frac{dy}{y}=\frac{1-x}{x^2}dx,""\int\frac{dy}{y}=\int\frac{1-x}{x^2}dx,""\int\frac{dy}{y}=\int\frac{1}{x^2}dx-\int\frac{dx}{x},""\ln y=-\frac{1}{x}-\ln x+C,"

where C is some constant.

If "y(1)=1," then "\ln 1=-1-\ln1+C."

From here C=1.

We will have

"\ln y=\ln e^{-\frac{1}{x}}-\ln x+\ln e," or"y=\frac{e^{1-\frac{1}{x}}}{x}."

Answer.

"y=\frac{e^{1-\frac{1}{x}}}{x}."